OpenStudy (yisselluna):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 1 y + 4 = 5x What is the resulting equation? 3x + 5x - 4 = 1 3x + 4 - 5x = 1 3x + 5x + 4 = 1

OpenStudy (natalief):

y+4=5x y=5x-4 Sub y=5x-4 into 3x+y=1, 3x + (5x-4) =1 3x+5x-4 =1